# Large Language Model-Brained GUI Agents: A Survey

### v3

@@ 1 Introduction
The rise of Large Language Models (LLMs) [^1] By LLMs, we refer to the general concept of foundation models capable of accepting various input modalities (e.g., visual language models (VLMs), multimodal LLMs (MLLMs)) while producing output exclusively in textual sequences [7]. [8, 9], especially those augmented with multimodal capabilities [10], has drastically redefined the possibilities for GUI automation. Beginning with models like ChatGPT [11], LLMs have demonstrated extraordinary proficiency in natural language understanding, code generation, and generalization across diverse tasks [12, 13, 8]. The integration of visual language models (VLMs) has further extended these capabilities, enabling these models to process visual data, such as the intricate layouts of GUIs [14]. This evolution bridges the gap between linguistic and visual comprehension, empowering intelligent agents to interact with GUIs in a more human-like and adaptive manner. By leveraging these advancements, LLMs and VLMs offer transformative potential, enabling agents to navigate complex digital environments, execute tasks dynamically, and revolutionize the field of GUI automation.
@@ 5.7.1 Computer Vision-Based GUI Parsing
Such CV-based GUI parsing layers provide essential information that greatly improves the agent’s ability to interact intuitively and accurately with diverse GUIs, especially in cases involving custom or non-standard elements that do not conform to typical accessibility protocols. This approach paves the way for more adaptable and resilient GUI agents, able to operate effectively in a wider variety of screen environments.
@@ 5.8 From Foundations to Innovations: A Roadmap
FoundationstoInnovations Frameworks Web [15, 185] [214, 215, 216, 217, 218, 219, 220, 221, 222, 223, 224, 225, 226, 227, 228, 229, 230, 231, 232, 233, 234] Mobile [16, 198, 145] [143, 235] [230, 231] [236, 237, 238, 239, 240, 241, 242, 243, 244, 245, 246] Computer [17, 148, 149] [198, 241, 231] [247, 248, 249, 246] Cross- Platform [230] [231, 198] [246, 250] Data Web [251, 252, 253, 254, 255, 256, 257] Mobile [169] [131, 235] [254, 256, 255] [258, 259, 260, 261, 262, 263, 264, 265, 266, 267, 268, 269, 270, 271, 272] Computer [169] [273, 271, 272] Cross- Platform [169, 271] [254, 256, 255, 257] [272] Models Foundation Models [79, 80, 150] [274, 275, 276, 277, 278, 279, 280, 281, 282, 283, 284] Web [23, 14, 169] [140, 229, 257] [285, 286, 287, 288, 289, 290, 291] Mobile [23, 14, 169] [212, 258, 268] [266, 292, 265, 293, 294, 295, 296, 146, 297, 298, 272, 289, 290, 291] Computer [23, 14, 273] [299, 298, 272] [169, 290] [291, 257] Cross- Platform [14, 23] [169, 272, 257] [298, 289, 290, 291] Evaluation Web [226, 218, 23] [129, 132, 133] [251, 252, 253, 254] [300, 301, 302, 303, 304, 305, 306, 307, 308, 309, 310, 311, 312, 313, 314, 315, 316, 317, 318] Mobile [269, 317] [319, 143, 131] [211, 145, 320, 321, 322, 262, 323, 324, 325, 326, 254, 23, 327, 328] Computer [23, 319] [317, 96, 329] [308, 309, 310] [330, 331, 332] Cross- Platform [248] [254, 309, 310] [319, 23, 317] Applications GUI Testing [333, 112] [334, 335, 336, 337, 338, 339, 340, 341, 342] Virtual Assistants [119] [343, 344, 345, 346, 347, 348, 349, 350, 351, 352, 353, 354, 355]
@@ 6.5 Takeaways
    Multi-Agent Synergy: Multi-agent systems, such as those in UFO and MMAC-Copilot, represent a significant trend in GUI agent development. By assigning specialized roles to different agents within a framework, multi-agent systems can enhance task efficiency, adaptability, and overall performance. As agents take on more complex tasks across diverse platforms, the coordinated use of multiple agents is proving to be a powerful approach, enabling agents to handle intricate workflows with greater precision and speed.
@@
    Multimodal Input Benefits: While some agents still rely solely on text-based inputs (e.g., DOM structures or HTML), incorporating visual inputs, such as screenshots, has shown clear performance advantages. Agents like WebVoyager and SeeAct highlight how visual data, combined with textual inputs, provides a richer representation of the environment state, helping agents make better-informed decisions. This integration of multimodal inputs is essential for accurate interpretation in visually complex or dynamic environments where text alone may not capture all necessary context.
@@
    Expanding Action Sets Beyond UI Operations: Recent agents have expanded their action sets beyond standard UI operations to include API calls and AI-driven actions, as seen in Hybrid Agent and AutoWebGLM. Incorporating diverse actions allows agents to achieve higher levels of interaction and task completion, particularly in environments where data can be directly retrieved or manipulated through API calls. This flexibility enhances agent capabilities, making them more efficient and adaptable across a wider range of applications.
@@
    Emerging Techniques for Improved Decision-Making: Novel approaches such as world models in WMA and search-based strategies in Search-Agent represent promising directions for more advanced decision-making. World models allow agents to simulate action outcomes, reducing unnecessary interactions and improving efficiency, especially in long-horizon tasks. Similarly, search-based algorithms like best-first and MCTS help agents explore action pathways more effectively, enhancing their adaptability in complex, real-time environments.
@@
    Toward Cross-Platform Generalization: Cross-platform frameworks, such as AutoGLM and OSCAR, underscore the value of generalizability in GUI agent design. These agents are pioneering efforts to create solutions that work seamlessly across mobile, desktop, and web platforms, moving closer to the goal of a one-stop GUI agent that can operate across multiple ecosystems. Cross-platform flexibility will be crucial for agents that aim to assist users consistently across their digital interactions.
@@ 7.4 Computer Agent Data
Although the desktop domain has fewer datasets compared to mobile and web, efforts like ScreenAgent highlight the growing interest and potential for developing sophisticated GUI agents for computer systems.
@@ 8.6 Cross-Platform Large Action Models
CogAgent [14] stands out as an advanced visual language model specializing in GUI understanding and navigation across PC, web, and Android platforms. Built on CogVLM [279], it incorporates a novel high-resolution cross-module to process GUI screenshots efficiently, enabling detailed comprehension of GUI elements and their spatial relationships. Excelling in tasks requiring OCR and GUI grounding, CogAgent achieves state-of-the-art performance on benchmarks like Mind2Web [251] and AITW [262]. Its ability to generate accurate action plans and interface with GUIs positions it as a pivotal step in developing intelligent agents optimized for GUI environments.
@@ 9.3 Evaluation Platforms
WebCanvas [304] Web 2024 Provides intermediate action tracking for realistic task assessment, along with an updated Mind2Web-Live dataset and tools for annotation. 542 tasks Step Success Rate, Task Success Rate, Efficiency Score Element Match, Text Match, trajectory length https://huggingface.co/datasets/iMeanAI/Mind2Web-Live
@@
Mind2Web-Live [226] Web 2024 Frames web navigation as functionality-driven question-answering, allowing agents to generalize across diverse web tasks and optimize navigation efficiency. 104 samples Task Success Rate, Efficiency Score Text Match, Image Match, Element Match, Path Length https://anonymous.4open.science/r/naviqate

### v4

@@ 1 Introduction
The rise of Large Language Models (LLMs) [^1] By LLMs, we refer to the general concept of foundation models capable of accepting various input modalities (e.g., visual language models (VLMs), multimodal LLMs (MLLMs)) while producing output exclusively in textual sequences [7]. [8, 9], especially those augmented with multimodal capabilities [10], has emerged as a game changer for GUI automation, redefining the the way agents interact with graphical user interfaces. Beginning with models like ChatGPT [11], LLMs have demonstrated extraordinary proficiency in natural language understanding, code generation, and generalization across diverse tasks [12, 13, 14, 8]. The integration of visual language models (VLMs) has further extended these capabilities, enabling these models to process visual data, such as the intricate layouts of GUIs [15]. This evolution bridges the gap between linguistic and visual comprehension, empowering intelligent agents to interact with GUIs in a more human-like and adaptive manner. By leveraging these advancements, LLMs and VLMs offer transformative potential, enabling agents to navigate complex digital environments, execute tasks dynamically, and revolutionize the field of GUI automation.
@@ 5.7.1 Computer Vision-Based GUI Parsing
Such CV-based GUI parsing layers provide critical grounding information that significantly enhances an agent’s ability to interact accurately and intuitively with diverse GUIs. This is particularly beneficial for handling custom or non-standard elements that deviate from typical accessibility protocols. Additionally, prompting methods like iterative narrowing have shown promise in improving the widget grounding capabilities of VLMs [197]. Together, these approaches pave the way for more adaptable and resilient GUI agents, capable of operating effectively across a broader range of screen environments and application contexts.
@@ 5.8 From Foundations to Innovations: A Roadmap
FoundationstoInnovations Frameworks Web [16, 186] [216, 217, 218, 219, 220, 221, 222, 223, 224, 225, 226, 227, 228, 229, 230, 231, 232, 233, 234, 235, 236, 237, 238, 239] Mobile [17, 200, 146] [144, 240, 238] [232, 233, 239] [241, 242, 243, 244, 245, 246, 247, 248, 249, 250, 251, 252] Computer [18, 149, 150] [200, 246, 233] [253, 254, 255, 251] [238, 239] Cross- Platform [232, 238] [233, 200] [251, 256] [239] Data Web [257, 258, 259, 260, 261, 262, 263, 264, 265, 266] Mobile [170, 132] [240, 260] [263, 261] [267, 268, 269, 270, 271, 272, 273, 274, 275, 276, 277, 278, 279, 280, 281, 266] Computer [170, 266] [282, 280, 281, 283] Cross- Platform [280, 266] [170, 263] [261, 264] [260, 281] Models Foundation Models [80, 81, 151] [284, 285, 286, 287, 288, 289, 290, 291, 292, 293, 294] Web [24, 15, 170] [141, 231, 264] [266, 295, 296, 297, 298, 299, 300, 301] Mobile [24, 15, 170] [214, 267, 277] [275, 302, 274, 303, 304, 305, 306, 147, 307, 308, 281, 299, 300, 301] [266] Computer [24, 15, 282] [309, 308, 281] [170, 300, 266] [301, 264, 283] Cross- Platform [15, 24, 266] [170, 281, 264] [308, 299, 300, 301] Evaluation Web [228, 220, 24] [130, 133, 134] [257, 258, 259, 260] [310, 311, 312, 313, 262, 314, 315, 316, 317, 318, 319, 320, 321, 322, 323, 324, 325, 326, 327] Mobile [278, 326] [328, 144, 132] [213, 146, 329, 330, 331, 271, 332, 333, 334, 335, 260, 24, 336, 337] Computer [24, 328] [326, 97, 338] [317, 318, 319] [339, 340, 341] Cross- Platform [254] [260, 318, 319] [328, 24, 326] Applications GUI Testing [342, 113] [343, 344, 345, 346, 347, 348, 349, 350, 351, 352] Virtual Assistants [120] [353, 354, 355, 356, 357, 358, 359, 360, 361, 362, 363, 364, 365]
@@
ScribeAgent [237] Web HTML-DOM Standard UI operations Single-agent architecture Specialized fine-tuning approach using production-scale workflow data to outperform general-purpose LLMs like GPT-4 in web navigation tasks https://github.com/colonylabs/ScribeAgent
@@
MobileGPT [252] Android mobile devices Simplified HTML representation Standard UI operations and navigation actions GPT-4-turbo for screen understanding and reasoning, GPT-3.5-turbo for slot-filling sub-task parameters Single-agent architecture augmented by a hierarchical memory structure Introduces a human-like app memory that allows for task decomposition into modular sub-tasks https://mobile-gpt.github.io
@@
AGUVIS [238] Web, desktop, and mobile Image-based observations Standard UI operations Fine-tuned Qwen2-VL [286] Single-agent architecture Pure vision-based approach for GUI interaction, bypassing textual UI representations and enabling robust cross-platform generalization https://aguvis-project.github.io
Ponder & Press [239] Web, Android, iOS Mobile, Windows, and macOS Purely visual inputs Standard UI operations GPT-4o and Claude 3.5 Sonnet for high-level task decomposition, a fine-tuned Qwen2-VL-Instruct [286] for GUI element grounding Divide-and-conquer architecture Purely vision-based GUI agent that does not require non-visual inputs https://invinciblewyq.github.io/ponder-press-page/
@@ 6.2 Mobile GUI Agents
MobileGPT [252] automates tasks on Android devices using a human-like app memory system that emulates the cognitive process of task decomposition—Explore, Select, Derive, and Recall. This approach results in highly efficient and accurate task automation. Its hierarchical memory structure supports modular, reusable, and adaptable tasks and sub-tasks across diverse contexts. MobileGPT demonstrates superior performance over state-of-the-art systems in task success rates, cost efficiency, and adaptability, highlighting its potential for advancing mobile task automation.
@@ 6.4 Cross-Platform GUI Agents
AGUVIS [238] leverages a pure vision approach to automate GUI interactions, overcoming limitations of text-based systems like HTML or accessibility trees. Its platform-agnostic design supports web, desktop, and mobile applications while reducing inference costs. AGUVIS employs a two-stage training process: the first focuses on GUI grounding, and the second integrates planning and reasoning within a unified model. This approach delivers state-of-the-art performance in both offline and online scenarios, streamlining decision-making and execution.
@@ 6.5 Takeaways
    Multi-Agent Synergy: Multi-agent systems, such as those in UFO [18] and MMAC-Copilot [200], represent a significant trend in GUI agent development. By assigning specialized roles to different agents within a framework, multi-agent systems can enhance task efficiency, adaptability, and overall performance. As agents take on more complex tasks across diverse platforms, the coordinated use of multiple agents is proving to be a powerful approach, enabling agents to handle intricate workflows with greater precision and speed.
@@
    Multimodal Input Benefits: While some agents still rely solely on text-based inputs (e.g., DOM structures or HTML), incorporating visual inputs, such as screenshots, has shown clear performance advantages. Agents like WebVoyager [219] and SeeAct [16] highlight how visual data, combined with textual inputs, provides a richer representation of the environment state, helping agents make better-informed decisions. This integration of multimodal inputs is essential for accurate interpretation in visually complex or dynamic environments where text alone may not capture all necessary context.
@@
    Expanding Action Sets Beyond UI Operations: Recent agents have expanded their action sets beyond standard UI operations to include API calls and AI-driven actions, as seen in Hybrid Agent [186] and AutoWebGLM [220]. Incorporating diverse actions allows agents to achieve higher levels of interaction and task completion, particularly in environments where data can be directly retrieved or manipulated through API calls. This flexibility enhances agent capabilities, making them more efficient and adaptable across a wider range of applications.
@@
    Emerging Techniques for Improved Decision-Making: Novel approaches such as world models in WMA [216] and search-based strategies in Search-Agent [224] represent promising directions for more advanced decision-making. World models allow agents to simulate action outcomes, reducing unnecessary interactions and improving efficiency, especially in long-horizon tasks. Similarly, search-based algorithms like best-first and MCTS help agents explore action pathways more effectively, enhancing their adaptability in complex, real-time environments.
@@
    Toward Cross-Platform Generalization: Cross-platform frameworks, such as AutoGLM [232] and OSCAR [251], underscore the value of generalizability in GUI agent design. These agents are pioneering efforts to create solutions that work seamlessly across mobile, desktop, and web platforms, moving closer to the goal of a one-stop GUI agent that can operate across multiple ecosystems. Cross-platform flexibility will be crucial for agents that aim to assist users consistently across their digital interactions.
  6. 6.
    Pure Vision-Based Agent: To enable universal GUI control, pure vision-based frameworks have emerged as a prominent solution. These agents rely solely on screenshots for decision-making, eliminating the need for access to metadata such as widget trees or element properties. Notable work like AGUVIS [238] exemplifies this approach. While pure vision-based methods offer greater generalizability and bypass system API limitations, they require strong “grounding” capabilities to accurately locate and interact with UI elements—an ability often lacking in many foundational models. Fine-tuning models specifically for visual grounding and GUI understanding, or integrating GUI parsing techniques like OmniParser [170], can address this challenge and enhance the agent’s ability to perform precise interactions.
@@ 7.1.2 Collection Pipeline
Mind2Web-Live [262] Web Sampled and re-annotated from the Mind2Web [257] Textual task descriptions, intermediate evaluation states, action sequences, and metadata, GUI screenshots 542 tasks, with 4,550 detailed annotation steps. Annotated by human experts. Emphasis on dynamic evaluation using “key nodes”, which represent critical intermediate states in web tasks. https://huggingface.co/datasets/iMeanAI/Mind2Web-Live
@@
AgentTrek [265] Web Web tutorials Task metadata, step-by-step instructions, action sequences, visual observations, reproducible native traces 4,902 trajectories VLM agent guided by tutorials, with Playwright capturing the traces Synthesizes high-quality trajectory data by leveraging web tutorials /
@@
LAM [283] Windows OS Application documentation, WikiHow articles, Bing search queries Task descriptions in natural language, step-by-step plans, action sequences, GUI screenshots 76,672 task-plan pairs, 2,192 task-action trajectories Instantiated using GPT-4, with actions tested and validated in the Windows environment using UFO [18] Provides a structured pipeline for collecting, validating, and augmenting data, enabling high-quality training for action-oriented AI models. https://github.com/microsoft/UFO/tree/main/dataflow
@@
Insight-UI [266] iOS, Android, Windows, Linux, Web Common Crawl corpus Textual tasks, plans, action sequences, GUI screenshots 434,000 episodes, 1,456,000 images Automatic simulations performed by a browser API Instruction-free paradigm and entirely auto-generated /
@@ 7.4 Computer Agent Data
The LAM dataset [283] is specifically designed to train and evaluate Large Action Models (LAMs) for GUI environments, bridging natural language task understanding and action execution. It comprises two core components: Task-Plan data, detailing user tasks with step-by-step plans, and Task-Action data, translating these plans into executable GUI actions. Sourced from application documentation, WikiHow articles, and Bing search queries, the dataset is enriched and structured using GPT-4. Targeting the Windows OS, with a focus on automating tasks in Microsoft Word, it includes 76,672 task-plan pairs and 2,688 task-action trajectories, making it one of the largest collections for GUI-based action learning. Data quality is ensured through a robust validation pipeline that combines LLM-based instantiation, GUI interaction testing, and manual review. Each entry is complemented with GUI screenshots and metadata, enabling models to learn both high-level task planning and low-level execution. The dataset’s modular design supports fine-tuning for specific GUI tasks and serves as a replicable framework for building datasets in other environments, marking a significant contribution to advancing GUI-based automation.
Although the desktop domain has fewer datasets compared to mobile and web, efforts like ScreenAgent and LAMs highlight the growing interest and potential for developing sophisticated GUI agents for computer systems.
@@ 8.2 Large Action Models
LAM [283] Windows OS Mistral-7B [369] 7B Task requests in natural language, application environmental data Plans, actions 76,672 task-plan pairs, 2,192 task-action trajectories The LAM model bridges the gap between planning and action execution in GUI environments. It introduces a multi-phase training pipeline combining task planning, imitation learning, self-boosting exploration, and reward-based optimization for robust action-oriented performance. https://github.com/microsoft/UFO/tree/main/dataflow
@@
Falcon-UI [266] iOS, Android, Windows, Linux, Web Qwen2-VL-7B 7B Screenshots of GUI with node information and OCR annotations for visible elements GUI actions and coordinates or bounding boxes for interaction elements Insight-UI dataset, further fine-tuned on datasets such as AITW, AITZ, Android Control, and Mind2Web Decouples GUI context comprehension from instruction-following tasks, leveraging an instruction-free pretraining approach. /
@@ 8.5 LAMs for Computer GUI Agents
Wang et al., [283] present a comprehensive overview of LAMs, a new paradigm in AI designed to perform tangible actions in GUI environments, using UFO [18] at Windows OS as a case study platform. Built on the Mistral-7B [369] foundation, LAMs advance beyond traditional LLMs by integrating task planning with actionable outputs. Leveraging structured inputs from tools like the UI Automation (UIA) API, LAMs generate executable steps for dynamic planning and adaptive responses. A multi-phase training strategy—encompassing task-plan pretraining, imitation learning, self-boosting exploration, and reinforcement learning—ensures robustness and accuracy. Evaluations on real-world GUI tasks highlight LAMs’ superior task success rates compared to standard models. This innovation establishes a foundation for intelligent GUI agents capable of transforming user requests into real-world actions, driving significant progress in productivity and automation.
@@ 8.6 Cross-Platform Large Action Models
CogAgent [15] stands out as an advanced visual language model specializing in GUI understanding and navigation across PC, web, and Android platforms. Built on CogVLM [289], it incorporates a novel high-resolution cross-module to process GUI screenshots efficiently, enabling detailed comprehension of GUI elements and their spatial relationships. Excelling in tasks requiring OCR and GUI grounding, CogAgent achieves state-of-the-art performance on benchmarks like Mind2Web [257] and AITW [271]. Its ability to generate accurate action plans and interface with GUIs positions it as a pivotal step in developing intelligent agents optimized for GUI environments. CogAgent has further evolved into its beta version, GLM-PC [398], offering enhanced control capabilities.
@@ 9.3 Evaluation Platforms
Mind2Web-Live [262] Web 2024 Provides intermediate action tracking for realistic task assessment, along with an updated Mind2Web-Live dataset and tools for annotation. 542 tasks Step Success Rate, Task Success Rate, Efficiency Score Element Match, Text Match, trajectory length https://huggingface.co/datasets/iMeanAI/Mind2Web-Live
Mind2Web-Live-Abstracted [228] Web 2024 Abstract the descriptions by omitting task-specific details and user input information in Mind2Web-Live, which are more streamlined and less time-consuming to compose. 104 samples Task Success Rate, Efficiency Score Text Match, Image Match, Element Match, Path Length https://anonymous.4open.science/r/naviqate
@@ 10.1 GUI Testing
LLMigrate [352] General testing Mobile Android GPT-4o DOM and screenshots Standard UI operations Automates the transfer of usage-based UI tests between Android apps Leverages multimodal LLMs to perform UI test transfers without requiring source code access /